OpenStudy (phenix):

There are 9 red marbles and 3 black marbles in a bag. Marbles are selected one at a time, with replacement. Each time, the color of the marble is recorded. Find the probability of selecting a red marble exactly 8 times in 10 selections.

OpenStudy (perl):

since you are replacing the marble, the probability of selecting a red marble stays the same, each time you draw from the bag.

OpenStudy (phenix):

So would that be 9/12 is the probability or

OpenStudy (perl):

Let's label choosing a red marble "S" for success. and choosing a non red marble "F" for fail. P(S) = 9/12 P(F) = 3/12 and you want probability of P( SSSSSSSSFF )

OpenStudy (perl):

the event "SSSSSSSSFF" means you drew a red marble on the first selection, the second selection, etc, and then on the ninth selection you drew a black marble and tenth you drew a black marble

OpenStudy (phenix):

alright, I understand that, so what's next?

OpenStudy (perl):

P(SSSSSSSSFF) = P(S) * P(S) * P(S) * ... P(F) * P(F) = 9/12 * 9/12 * .. 3/12 * 3/12 = (9/12) ^8 * (3/12) ^2

OpenStudy (perl):

however, we are not done, because there are more ways to select 8 red marbles and 2 black marbles. you could select 2 black marbles first, then 8 red marbles. : FF SSSS SSSS you could select a red marble, then a black marble, then a red marble...: S F S F SSSS SS

OpenStudy (perl):

but each arrangement of 8 S's and 2 F's has the same probability of (9/12)^8 * (3/12)^2

OpenStudy (phenix):

so that is the answer itself, (9/12)^8 * (3/12)^2?

OpenStudy (perl):

its part of it

OpenStudy (perl):

P( exactly eight S two F ) = (# ways to rearrange 8 S's,2 F's) * (9/12)^8 * (3/12)^2

OpenStudy (phenix):

ohhhh okay so that is all?

OpenStudy (perl):

right, so you have to find the number of ways to rearrange SSSS SSSS FF. now doing it by hand takes too long. we have a formula for it. 10! / ( 8! 2! )

OpenStudy (phenix):

Oh jeez haha so how do i plug that into the formula?

OpenStudy (perl):

$$ \Large P( exactly~ 8 ~ red~ marbles ~ 2~black) \\\Large = \frac{10!}{8! ~2!} \cdot \left( \frac{9}{12} \right)^8 \cdot \left(\frac{3}{12}\right)^2 $$

OpenStudy (phenix):

What does the "!" mean in that equation?

OpenStudy (perl):

8! = 8x7x6x5x4x3x2x1 2! = 2x1

OpenStudy (perl):

you can also use the expression 10 choose 8. some calculators allow you to do that nCr on TI 83

OpenStudy (phenix):

Is the answer 5/254803968?

OpenStudy (alexandervonhumboldt2):

simple 1-10 ! table 1!=1 2!=2 3!=6 4!=24 5!=120 6!=720 7!=5040 8!=40320 9!=362880 10!=3628800

OpenStudy (alexandervonhumboldt2):

! means a factorial a factorial of a numebr is a multiplication of all previous numebr exept 0 that are natural including the numebr

OpenStudy (perl):

$$ \Large { \frac{10!}{8! ~2!}\\~\\ = \frac{10 \times 9\times 8\times 7\times 6\times 5\times 4 \times 3\times 2\times 1 }{(8\times 7\times 6\times 5\times 4 \times 3\times 2\times 1)\cdot (2 \times 1 ) } } $$

OpenStudy (perl):

but you can simplify this, notice you can cancel 8's, 7's, , etc

OpenStudy (phenix):

Alright I get that but I don't have a calculator on me I'm using my phone so it's not a graphic calc so i'm not sure if it calculates probability

OpenStudy (perl):

$$ \Large { \frac{10!}{8! ~2!}\\~\\ = \frac{10 \times 9\times 8\times 7\times 6\times 5\times 4 \times 3\times 2\times 1 }{(8\times 7\times 6\times 5\times 4 \times 3\times 2\times 1)\cdot (2 \times 1 ) } \\~\\ = \frac{10 \times 9\times \color{red}{\not 8}\times \color{blue}{\not7}\times \color{violet}{\not 6}\times 5\times 4 \times 3\times 2\times 1 }{(\not \color{red}{\not 8}\times \color{blue}{\not7}\times \color{violet}{\not 6}\times 5\times 4 \times 3\times 2\times 1)\cdot (2 \times 1 ) } } $$

OpenStudy (phenix):

and so it cancels out until its 10x9/2?

OpenStudy (perl):

$$ \Large { \frac{10!}{8! ~2!}\\~\\ = \frac{10 \times 9\times 8\times 7\times 6\times 5\times 4 \times 3\times 2\times 1 }{(8\times 7\times 6\times 5\times 4 \times 3\times 2\times 1)\cdot (2 \times 1 ) } \\~\\ = \frac{10 \times 9\times \color{red}{\not 8}\times \color{blue}{\not7}\times \color{violet}{\not 6}\times 5\times 4 \times 3\times 2\times 1 }{(\not \color{red}{\not 8}\times \color{blue}{\not7}\times \color{violet}{\not 6}\times 5\times 4 \times 3\times 2\times 1)\cdot (2 \times 1 ) } \\~\\ \\ \therefore \\ = \frac{10\times 9 }{2\times 1} } $$

OpenStudy (perl):

$$ \Large{ P( exactly~ 8 ~ red~ marbles ~ 2~black) \\\Large = \frac{10!}{8! ~2!} \cdot \left( \frac{9}{12} \right)^8 \cdot \left(\frac{3}{12}\right)^2 \\ ~\\= \frac{10\times 9 }{2\times 1}\cdot \left( \frac{9}{12} \right)^8 \cdot \left(\frac{3}{12}\right)^2 \\~\\ = \frac{90}{2} \cdot \left( \frac{9}{12} \right)^8 \cdot \left(\frac{3}{12}\right)^2 \\~\\ = 45 \cdot \left( \frac{9}{12} \right)^8 \cdot \left(\frac{3}{12}\right)^2 }$$

OpenStudy (phenix):

So the answer is .28156? Or are there more steps?

OpenStudy (perl):

you're done :P

OpenStudy (phenix):

thank you so much!!!!!!!!!!!!!!!!